OpenStudy (anonymous):

40 circuits is 80% of what number of circuits?

OpenStudy (anonymous):

40*0.80

OpenStudy (anonymous):

If you would multiply 40 to .80 since it's a percent, you should get 32 circuits.

OpenStudy (nincompoop):

40/x = 80/100

OpenStudy (nincompoop):

solve for x x/40 = 100/80 x = 100/80 * 40